# Efficient sequential Bayesian inference for state-space epidemic models using ensemble data assimilation

### v1

@@ Appendix F Additional results on the mpox dynamics
This section presents supplementary figures that support the results discussed in the main text. Figure F.1 illustrates the application to the weekly mpox data using a Poisson observation model. Figure F.2 shows the filtered trajectories of the unobserved SEIR compartments, and Figure F.3 presents the estimated parameters obtained with the SMC2 and eSMC2 algorithms.
Figure F.1: Inference of daily incidence and effective reproduction number for the 2022 U.S. monkeypox outbreak under Poisson observation variance.
Figure F.2: Unobserved states of the SEIR model for the mopox data under the Negative Binomial observation model. Filtering mean and 95% credible interval are shown for $S_{t},\penalty 10000\ E_{t},\penalty 10000\ I_{t},$ and $\beta_{t}$.
Figure F.3: Comparison of the filtered estimate of the parameters for the mpox data. The graph shows the filtered means and corresponding 95% credible interval

### v2

@@ Appendix F Additional results on the mpox dynamics
This section presents supplementary results supporting the analyses in the main text. We first illustrate inference under a Poisson observation model, then present additional results under a Negative Binomial observation model, and finally discuss forecasting performance comparing eSMC2 with a simpler autoregressive baseline.
Inference under Poisson observation variance
Figure F.1 shows estimated weekly incidence and the effective reproduction number for the 2022 U.S. mpox outbreak assuming a Poisson observation model. The filtering procedure captures the overall epidemic trajectory under this simplified variance assumption.
Figure F.1: Inference of weekly incidence and effective reproduction number under Poisson observation variance. Filtering mean and 95% credible intervals are shown for reported mpox incidence and estimated $R_{t}$.
Additional results under Negative Binomial observation variance
Figures F.2 and F.3 present results obtained under a Negative Binomial observation model.
Figure F.2: Unobserved SEIR states under Negative Binomial observation variance. Filtering mean and 95% credible intervals are shown for $S_{t},E_{t},I_{t},$ and $\beta_{t}$.
Figure F.3: Comparison of filtered parameter estimates under Negative Binomial observation variance. Filtered means and 95% credible intervals obtained with SMC2 and eSMC2 are shown.
Autoregressive baseline model for comparison
For reference, we implemented a discrete-time autoregressive model of order $p$:
$y_{t}=a+\sum_{k=1}^{p}b_{k}y_{t-k}+\varepsilon_{t},$
where $Y_{t}$ is the observed incidence at time $t$, $a$ is an intercept, $b_{k}$ are autoregressive coefficients, and $\varepsilon_{t}\sim\mathcal{N}(0,\sigma^{2})$ represents the innovation term. Predictive distributions were generated by iteratively simulating forward trajectories conditional on the last $p$ observed values. This baseline captures short-term temporal dependence but does not explicitly model latent transmission dynamics.
Forecast evaluation metrics
Let $F_{{}_{T+h}}$ denote the predictive distribution at horizon T+h and $y_{{}_{T+h}}$ the observed incidence. Forecast performance was evaluated using the following metrics:
  • Mean Absolute Error (MAE):
    $\text{MAE}=\frac{1}{H}\sum_{h=1}^{H}|y_{{}_{T+h}}-\hat{m}_{{}_{T+h}}|,$
    where $H$ is the forecast horizon and $\hat{m}_{{}_{T+h}}$ is the predictive median.
  • Weighted Interval Score (WIS): To jointly assess probabilistic calibration and sharpness, we compute WIS using central $(1-\alpha)$ prediction intervals (Bracher et al., 2021). Let $l_{\alpha,_{T+h}}$ and $u_{\alpha,_{T+h}}$ be the lower and upper bounds of the $(1-\alpha)$ interval from $F_{{}_{T+h}}$. The interval score at horizon T+h is
    $\text{IS}_{\alpha}(F_{{}_{T+h}},y_{{}_{T+h}})=(u_{\alpha,_{T+h}}-l_{\alpha,_{T+h}})+\frac{2}{\alpha}(l_{\alpha,_{T+h}}-y_{{}_{T+h}})\mathbf{1}\{y_{{}_{T+h}}<l_{\alpha,_{T+h}}\}+\frac{2}{\alpha}(y_{{}_{T+h}}-u_{\alpha,_{T+h}})\mathbf{1}\{y_{{}_{T+h}}>u_{\alpha,_{T+h}}\}$
    The WIS combines $K$ intervals and the absolute error of the median:
    $\text{WIS}=\frac{1}{H}\sum_{h=1}^{H}\left[\frac{1}{K+1/2}\left(w_{0}|y_{{}_{T+h}}-\hat{m}_{{}_{T+h}}|+\sum_{k=1}^{K}w_{k}\text{IS}_{\alpha_{k}}(F_{{}_{T+h}},y_{{}_{T+h}})\right)\right],$
    where $w_{0}=1/2$ and $w_{k}=\alpha_{k}/2$ are the weights of individual interval scores.
  • Empirical 95% prediction interval coverage: To measure the proportion of observed outcomes that fall within the model’s 95% predictive interval, defined as
    $\text{Coverage}_{95}=\frac{1}{H}\sum_{h=1}^{H}\mathbf{1}\{y_{{}_{T+h}}\in[l_{0.95,_{T+h}},u_{0.95,_{T+h}}]\},$
    where $l_{0.95,_{T+h}}$ and $u_{0.95,_{T+h}}$ are the lower and upper bounds of the 95% prediction interval at horizon T+h.
Long-term forecast performance using eSMC2
Figure F.4 presents forecasts for the final outbreak phase. The model captures the overall declining trend, though short-term daily fluctuations are less well predicted. Predictions are well calibrated, with 95% posterior credible intervals covering all out-of-sample observations. These results confirm that eSMC2 produces reliable probabilistic forecasts for real-time epidemic monitoring.
Figure F.4: Long-term forecast performance with eSMC2. Forecast distributions with 50%, 75%, 90%, and 95% credible intervals are shown in purple. In-sample observations (red) occur before the forecast start date (vertical dashed line), while out-of-sample observations (yellow) are used for validation.